# Distal Causal Excursion Effects: Modeling Long-Term Effects of Time-Varying Treatments in Micro-Randomized Trials

### v1

@@ 4.2 Proposed Estimator and Comparator Methods
We consider the proposed estimators $\hat{\beta}$ (the one without cross-fitting) and $\tilde{\beta}$ (with $K=5$ fold cross-fitting). In both implementations, the working models for the nuisance parameters $\mu_{t}(H_{t},1)$ and $\mu_{t}(H_{t},0)$ are pooled over $t\in[T]$ and are fitted using the generalized additive model (gam in R package mgcv). Each working model includes covariates $X_{t}$ and $Z_{t}$, where the continuous covariate $X_{t}$ is included as a penalized spline with cubic spline basis and the binary covariate $Z_{t}$ is included untransformed. These working models are misspecified because of omitted variables ($\mu_{t}$ depends on all the history variables in addition to $X_{t},Z_{t}$) and incorrect functional form (the dependence of $\mu_{t}$ on $X_{t}$ and $Z_{t}$ varies with $t$, which is not captured in the working model that pools over all $t\in[T]$). We introduce this misspecification to illustrate the robustness of the proposed estimator to model specificiation in $\hat{\mu}$ (Remark 4).
@@
The DCEE is a novel estimand and no existing method in the literature is directly applicable. Nonetheless, we choose as comparators two approaches commonly used to analyze longitudinal data with time-varying treatments to illustrate the advantage of the proposed method in estimating the DCEE: the generalized estimating equations (GEE, Liang & Zeger, 1986) and the structural nested mean model (SNMM, Robins, 1994). To use GEE to estimate the marginal DCEE $\beta^{\star}_{1}$, we specify the mean model for $Y$ as a linear combination of $A_{t},X_{t},Z_{t}$, and we pool data over all $t\in[T]$ when fitting the model. The coefficient for $A_{t}$ (same across all $t$) is the GEE estimator for $\beta^{\star}_{1}$. To use GEE to estimate the moderated DCEE $\beta^{\star}_{2}$ and $\beta^{\star}_{3}$, we specify the mean model for $Y$ as a linear combination of $A_{t},X_{t},Z_{t},A_{t}\times Z_{t}$ and pool data over all $t\in[T]$ when fitting the model. The coefficients for $A_{t}$ and $A_{t}\times Z_{t}$ are the GEE estimators for $\beta^{\star}_{2}$ and $\beta^{\star}_{3}$, respectively. We use a working independence correlation matrix and the robust standard error estimator.
@@
For SNMM, we used the g-estimation-based implementation in R package DTRreg. To estimate the marginal DCEE $\beta^{\star}_{1}$, we specify the blip model for each $t\in[T]$ to only include an intercept. The treatment assignment model is correctly specified with $X_{t}$ and $Z_{t}$ included in the model for $A_{t}$, and because DTRreg fits separate models for each $t\in[T]$, the dependence of $A_{t}$ on $t$ is correctly modeled as well. The treatment-free model for mean outcome after the treatment blips at and after $t$ are removed is specified as a linear model that depends on $X_{t}$ and $Z_{t}$. This choice of the treatment-free model mimics the nuisance model for $\mu_{t}$ in the proposed method, so as to make the SNMM modeling as comparable to the proposed method as possible. To use SNMM to estimate the moderated DCEE $\beta^{\star}_{2}$ and $\beta^{\star}_{3}$, we add $Z_{t}$ into the blip model.
@@ 4.3 Results
We conducted simulations for sample sizes $n=30,50,100,200,300,400,500$, each with 1000 replicates. Results are shown in Fig. 2. The proposed $\hat{\beta}$ and $\tilde{\beta}$ have satisfactory performance for sample size as small as 30: for all estimands, bias is close to 0 and the coverage of 95% confidence intervals is close to the nominal level. There is no substantial difference between the implementations with and without cross-fitting. The two comparator methods, GEE and SNMM, suffer from substantial bias and poor coverage, because they were not designed for estimating the DCEE.
Figure 2: The numerical performance of DCEE, DCEE-cf, GEE, and SNMM for the three estimands in terms of bias (top panel), standard deviation (middle panel), and the coverage of 95% confidence intervals.
@@ 5 Application
We use data from HeartSteps I, the first of a sequence of HeartSteps MRTs, which aimed to design mHealth interventions that support sedentary adults in reaching and sustaining recommended physical activity levels (Klasnja et al., 2015). In this application, we focus on the micro-randomized activity suggestions, which are one of the interventions tested in the MRT. $n=37$ participants were enrolled in the study for six weeks. Decision points occurred five times per day at pre-determined times, with a total of $T=210$ decision points per person. At each decision point, a participant was eligible to be randomized if they were not driving or walking and their phone had a stable internet connection. At every eligible decision point, a participant was randomized with probability 0.6 to receive an activity suggestion, which is a push notification suggesting brief walking or stretching activity, and with probability 0.4 to receive no intervention. Across all participants, approximately 80% of the decision points turned out to be eligible. Each participant wore a wristband tracker throughout the study, which continuously recorded their step count.
@@
The long-term goal of these activity suggestions was to help participants develop a habit for being physically active. Therefore, in this illustration we consider as distal outcome the average daily step count in the last week of the study, a proxy for positive habit formation. We analyze the effect of the activity suggestions delivered in the first five weeks on this distal outcome measured in Week 6. We discuss the choice of this distal outcome at the end of this section.
@@
Throughout this section we set $\omega(t)=1/T$. In fitting the nuisance parameters $\mu_{t}(H_{t},1)$ and $\mu_{t}(H_{t},0)$, we use the generalized additive model (gam in R package mgcv), pool data over all $t\in[T]$ when fitting the model, and include the step count in the 30 minutes prior to decision point $t$ and a location indicator for whether the person is at home or work (location indicator $=1$) or elsewhere (location indicator $=0$).
@@
We first study the fully marginal excursion effect by setting $S_{t}=\emptyset$ and $f(t,s)=1$ in (6). The estimated effect is $\hat{\beta}=69$ (95% CI $=[-107,245]$). This indicates that when averaged over the intervention period under consideration (Weeks 1 through 5), sending an activity suggestion, compared to not sending one, does not have a detectable effect on the distal outcome measured in Week 6. We further investigate the effect moderation by the decision point index $t$, i.e., the length of time a participant has spent in the study, because prior studies have found a deteriorating effect over time on a short-term proximal outcome measured after each decision point (Klasnja et al., 2018). We set $S_{t}=\emptyset$, $f(t,s)=(1,t-1)^{T}$, and $\beta=(\beta_{0},\beta_{1})^{T}$. The transformation $t-1$ in $f(t,s)$ ensures the interpretation of $\beta_{0}$ as the effect on the distal outcome for the treatment at the first decision point. The result is $\hat{\beta}_{0}=344$ (95% CI $=[-74,761]$) and $\hat{\beta}_{1}=-3.4$ (95% CI $=[-7.5,0.7]$). This means that the activity suggestion at the beginning of the study possibly has a sizeable effect (increasing the average daily step count in Week 6 by 344 steps) but this effect likely deteriorates over time (due to a negative $\hat{\beta}_{1}$). This is an intriguing finding: The interventions delivered at the beginning of the study are more distant from the proxy for habit formation measured in Week 6, but their effects on habit formation are actually larger than the interventions sent later in the study.
@@
To investigate whether the linearity assumption made in the effect moderation analysis [where we set $f(t,s)=(1,t-1)^{T}$] is reasonable, we conducted another moderation analysis where $f(t,s)$ contains an intercept and B-spline basis of 6 degrees of freedom. This allows the causal excursion effect of $A_{t}$ on $Y$ [i.e., $\tau(t,s)$] to vary flexibly with $t$. To enhance the interpretability of the result, we constructed the B-spline basis on $\text{day}_{t}:=\lfloor\frac{t-1}{5}\rfloor+1$. Fig. 3 shows that there is an increasing trend in the DCEE during the first week of the study. After the effect reached its peak by the end of Week 1, it starts to gradually decrease and crosses 0 at around the end of Week 2.
@@
Figure 3: How the effect of a push notification on the distal outcome measured in Week 6 varies across different days of the study in the HeartSteps MRT. Solid curve represents the point estimate and shaded area represents 95% pointwise confidence intervals.
@@
This finding suggests that early-in-the-study activity suggestions in the HeartSteps study may play a crucial role in shaping long-term behavior change, even if they are more distant in time to the distal outcome. One possible explanation is that interventions delivered early in the study help establish an initial momentum for behavior change, laying the foundation for participants to internalize and sustain new activity patterns over time. These early nudges may have a lasting impact by fostering psychological mechanisms such as self-efficacy, intrinsic motivation, or goal-setting behaviors.
@@
In contrast, activity suggestions provided later in the study have diminishing returns, due to two possible explanations. First, some participants may have already developed a habit later in the study and thus rely less on external prompts. Second, some participants may have already disengaged due to message fatigue or perceived redundancy in the prompts. This aligns with theories of habit formation, which suggest that consistent reinforcement in the early stages is crucial for establishing automaticity, while interventions introduced too late may struggle to alter established behavioral patterns (Lally & Gardner, 2013).
This finding highlights the importance of timing in digital interventions. It suggests that future mHealth interventions should prioritize delivering more intensive or strategically timed support early in the intervention period to promote long-term habit formation.
@@
Lastly, we note that the activity suggestions were also randomly delivered in Week 6, and thus the distal outcome was also influenced by these interventions. Our analysis is still valid in the presence of such dependence, but the distal outcome and the analysis result should be viewed in the context of the Week 6 interventions. In other words, had the micro-randomizations been stopped by the end of Week 5 and the participants continued to wear the tracker in Week 6 allowing the distal outcome to still be measured, the DCEE in that hypothetical setting could be different due to potential interactions between treatments in Week 6 and those in earlier weeks.
@@ 6 Discussion
We introduced the distal causal excursion effect (DCEE), a novel causal estimand for evaluating the long-term impact of time-varying treatments in MRTs. The DCEE enables a parsimonious and interpretable model for the causal effect even when treatments are randomized at hundreds or thousands of decision points. We developed two robust estimators, one with cross-fitting and one without, both of which are robust to misspecification of the outcome model. Through asymptotic theory, simulations, and an application to the HeartSteps MRT, we demonstrated the feasibility and practical utility of our approach. For HeartSteps, our findings reveal that early-in-the-study activity suggestions may have a stronger long-term impact on habit formation at the end of the study. This work provides the critically needed toolkit for scientists working on digital interventions to assess long-term causal effects using MRT data.
In DCEE, we set the reference policy to be the MRT policy, and the excursion policies are one-step excursions that only differ from the reference policy at decision point $t$. The DCEE definition can be straightforwardly generalized to other reference policies and excursion policies. However, the feasibility of estimation may not easily generalize depending on how different these policies are from the MRT policy under which the data is collected, due to the potential instability of inverse probability weights. For estimating such effects, methods to stabilize the inverse weights could be explored as a future direction.
Future work could extend our approach in several directions. First, our method assumes no unmeasured confounding and a correctly specified propensity score, and both assumptions are satisfied in MRTs but not necessarily in observational studies; incorporating doubly robust estimation and methods to address unmeasured confounding would improve applicability to observational studies. Second, the current linear model for the DCEE could be extended to data-adaptive models, especially if the set of moderators $S_{t}$ is large. Third, the current approach estimates an average causal effect across all participants; future research could extend the method to estimate heterogeneous person-specific treatment effects. Fourth, linking proximal and distal outcomes through mediation analysis could clarify the mechanisms driving long-term behavior change. Finally, future work could explore policy learning approaches to identify the optimal sequence of treatments that yield the greatest long-term benefit.
@@
Supplementary Material A contains the extensions and details for the DCEE examples in Section 2.2. Supplementary Material B proves the identification result Theorem 1. Supplementary Material C proves the weighted average form of $\beta^{\star}$ given in Remark 3. Supplementary Material D derives the estimating function $\phi(\beta,\mu)$ in Eq. 10. Supplementary Material E proves the asymptotic normality Theorem 2. Supplementary Material F details the numerical computation of true parameter values in simulation studies.
@@
We provide the detailed derivation for the four DCEE Examples in Section 2.2, where because Assumption 1 holds in all these examples, we can invoke Theorem 1 and in particular Eq. 5 to calculate $\tau(t)$. We also present more general versions of Example 3 and Example 4 to illustrate additional nuanced effects that are captured by the DCEE.

### v2

@@ 4.2 Proposed Estimator and Comparator Methods
We consider the proposed estimators $\widehat{\beta}$ (without cross-fitting) and $\widetilde{\beta}$ (with $K=5$ fold cross-fitting). In both estimators, the working models for nuisance parameters $\mu_{t}(H_{t},1)$ and $\mu_{t}(H_{t},0)$ are pooled over $t\in[T]$ and fitted using either the generalized additive model (gam in R package mgcv (Wood, 2011)) or a fast implementation of random forests (R package ranger (Wright & Ziegler, 2017)). Each working model includes covariates $X_{t}$ (as a penalized cubic spline in gam) and untransformed $Z_{t}$. These working models are misspecified due to omitted history variables (because $\mu_{t}$ depends on the full history in addition to $X_{t},Z_{t}$) and incorrect functional form (the dependence of $\mu_{t}$ on $X_{t}$ and $Z_{t}$ varies with $t$, which is not captured in the working model that pools over all $t\in[T]$). We introduce this misspecification to illustrate the robustness of the proposed estimator to model specificiation in $\widehat{\mu}$ (Remark 5).
@@
The DCEE is a novel estimand, with no existing method directly applicable. Nonetheless, we include as comparators three common approaches for analyzing longitudinal data with time-varying treatments, to illustrate the advantage of our method: the generalized estimating equations (GEE, Liang & Zeger, 1986), the structural nested mean model (SNMM, Robins, 1994), and the weighted and centered least squares (WCLS, Boruvka et al., 2018). To use GEE for estimating the marginal DCEE $\beta^{\star}_{1}$, we specify the mean model for $Y$ as a linear combination of $A_{t},X_{t},Z_{t}$, pooling data over all $t\in[T]$; the coefficient for $A_{t}$ (same across all $t$) is the GEE estimator for $\beta^{\star}_{1}$. For the moderated DCEE $\beta^{\star}_{2}$ and $\beta^{\star}_{3}$, we specify the mean model for $Y$ as a linear combination of $A_{t},X_{t},Z_{t},A_{t}\times Z_{t}$, again pooling over $t$. The coefficients for $A_{t}$ and $A_{t}\times Z_{t}$ are the GEE estimators for $\beta^{\star}_{2}$ and $\beta^{\star}_{3}$, respectively. We use a working independence correlation matrix and the robust standard error estimator.
@@
For SNMM, we used the g-estimation implementation in the R package DTRreg (Wallace et al., 2017). To estimate the marginal DCEE $\beta^{\star}_{1}$, we specified a blip model with intercept only for each $t\in[T]$. The treatment assignment model included $X_{t}$ and $Z_{t}$ and correctly captured time dependence because DTRreg fits separate models for each $t$. The treatment-free outcome model was specified as a linear model depending on $X_{t}$ and $Z_{t}$, mimicking our nuisance model for $\mu_{t}$ to ensure comparability. For estimating the moderated DCEE parameters $\beta^{\star}_{2}$ and $\beta^{\star}_{3}$, we additionally included $Z_{t}$ in the blip model.
For WCLS, we used the wcls function in R package MRTAnalysis. We set the proximal outcome argument as the distal outcome (same value repeated for each decision point), and used as control variables a linear form of $X_{t}$ and $Z_{t}$. The moderator argument is set to intercept only (for estimating $\beta^{\star}_{1}$) or include $Z_{t}$ (for estimating $\beta^{\star}_{2}$ and $\beta^{\star}_{3}$). The output proximal CEE are treated as the estimated DCEE in the simulation.
@@ 4.3 Results
We conducted simulations for sample sizes $n=30,50,100,200,300,400,500$, each with 1000 replicates. Results are shown in Fig. 2. The proposed DCEE estimators perform well even at $n=30$: for all estimands, bias is near zero and 95% confidence interval coverage is close to nominal. There is no substantial difference between versions with and without cross-fitting. In contrast, the comparator methods GEE, SNMM, and WCLS show substantial bias and poor coverage, because they were not designed for estimating the DCEE.
Figure 2: The numerical performance of DCEE (using GAM or random forests for nuisance parameter estimation, and with or without cross-fitting) GEE, SNMM, and WCLS for the three estimands in terms of bias (plots A–C), standard deviation (plots D–F), and the coverage of 95% confidence intervals (plots G–I).
We observed very similar finite-sample performance between the cross-fitted and non-cross-fitted versions of the proposed estimator when the nuisance outcome model was estimated using either GAM or random forests. This may reflect that the nuisance functions in our generative setting were not overly complex and could be reliably learned by these algorithms. Given that our estimating equation is globally robust and does not rely on the nuisance model for identification, both approaches perform similarly in this context. A more comprehensive investigation of their relative merits is an important direction for future work. In practice, we recommend using cross-fitting when employing highly flexible learners in more complex settings or when overfitting bias is a concern (Zivich & Breskin, 2021; Naimi et al., 2021), but using the non-cross-fitted version if cross-fitting fails to converge (e.g., due to very small sample sizes).
@@ 5 Application
We use data from HeartSteps I, the first in a sequence of HeartSteps MRTs, designed to support sedentary adults in achieving and sustaining recommended physical activity levels (Klasnja et al., 2015). In this application, we focus on the micro-randomized activity suggestions. $n=37$ participants were enrolled for 6 weeks, with decision points five times per day at pre-specified times, totaling $T=210$ per person. At each decision point, a participant was eligible to be randomized if not driving or walking and if their phone had a stable internet connection. At every eligible decision point, a participant was randomized with probability 0.6 to receive an activity suggestion (a push notification suggesting brief walking or stretching) and probability 0.4 to receive no intervention. Across all participants, approximately 80% of the decision points were eligible. Each participant wore a wristband tracker that continuously recorded step count.
@@
The long-term goal of these activity suggestions was to help participants develop a habit of being physically active. Therefore, in this illustration we use the average daily step count in Week 6 as the distal outcome, serving as a proxy for habit formation. We analyze the effect of suggestions delivered in the first 5 weeks on this distal outcome. We discuss the distal outcome choice at the end of this section.
@@
We set $\omega(t)=1/T$. To fit nuisance parameters $\mu_{t}(H_{t},1)$ and $\mu_{t}(H_{t},0)$, we use the generalized additive model, pooling data over all $t\in[T]$, and include the step count in the prior 30 minutes and a location indicator (home/work =1=1, elsewhere =0=0).
@@
We first examine the fully marginal excursion effect by setting $S_{t}=\emptyset$ and $f(t,s)=1$ in (6). The estimated effect is $\widehat{\beta}=69$ (95% CI =[107,245]=[-107,245]), which indicates that, averaged over Weeks 1–5, sending an activity suggestion has no detectable effect on Week 6 step count. We then assess the effect moderation by the decision point index $t$ (proportional to time in study), motivated by prior evidence of a diminishing effect over time on a short-term proximal outcome measured after each decision point (Klasnja et al., 2019). We set $S_{t}=\emptyset$, $f(t,s)=(1,t-1)^{T}$, and $\beta=(\beta_{0},\beta_{1})^{T}$, where the t1t-1 ensures that $\beta_{0}$ reflects the effect at the first decision point. We find $\widehat{\beta}_{0}=344$ (95% CI =[74,761]=[-74,761]) and $\widehat{\beta}_{1}=-3.4$ (95% CI =[7.5,0.7]=[-7.5,0.7]). This indicates that activity suggestions at the beginning of the study may have a sizeable effect (increasing Week 6 daily step count by 344 steps), but this effect likely deteriorates over time (due to the negative $\widehat{\beta}_{1}$). This is intriguing: early interventions, though more distant from Week 6, appear to have larger effects on habit formation than those delivered later.
@@
To assess the linearity assumption [where we set $f(t,s)=(1,t-1)^{T}$], we conducted another moderation analysis with $f(t,s)$ containing an intercept and a B-spline basis (6 degrees of freedom), allowing $\tau(t,s)$ to vary flexibly with $t$. For interpretability, we constructed the B-spline basis on $\text{day}_{t}:=\lfloor\frac{t-1}{5}\rfloor+1$. Fig. 3 shows an increasing trend in the DCEE during the first week, peaking at the end of Week 1, then gradually decreasing and crossing 0 at around the end of Week 2.
@@
Figure 3: Time-varying effect of a push notification on the distal outcome (Week 6 average daily steps) across study days in the HeartSteps MRT. Solid curve: point estimate; shaded area: 95% pointwise confidence intervals.
@@
This finding suggests that early activity suggestions in the HeartSteps study may play a crucial role in shaping long-term behavior change, even if they are more distant from the distal outcome. One possible explanation is that early interventions help establish a momentum for behavior change, laying a foundation for participants to internalize and sustain new activity patterns. These early nudges may have a lasting impact by fostering psychological mechanisms such as self-efficacy, intrinsic motivation, or goal-setting.
@@
In contrast, activity suggestions provided later in the study show diminishing returns, for two possible reasons that align with habit formation theory (Lally & Gardner, 2013). First, later in the study some participants may have already developed a habit and thus rely less on prompts. Second, some participants may have disengaged due to message fatigue or perceived redundancy in the prompts. Our finding suggests that future mHealth programs should prioritize timely support early on to promote long-term habit formation.
@@
Lastly, we note that the activity suggestions were also randomly delivered in Week 6, and thus the distal outcome was also influenced by those interventions. Our analysis remains valid despite of this dependence, but the results should be interpreted in the context of the Week 6 interventions. In other words, if micro-randomizations had stopped at the end of Week 5, the DCEE in that hypothetical scenario could be different due to potential interactions between Week 6 interventions and earlier interventions.
@@ 6 Discussion
We introduced the distal causal excursion effect (DCEE), a novel causal estimand for evaluating the long-term impact of time-varying treatments in MRTs. We developed two estimators, both robust to outcome model misspecification. When applied to HeartSteps MRT, our found that early-in-the-study activity suggestions may have a stronger long-term impact on habit formation. Overall, this work provides the much-needed toolkit for scientists developing digital interventions to assess long-term causal effects using MRT data.
The causal excursion effect (CEE) and DCEE offer valuable but distinct insights for optimizing future digital interventions. CEE provides information on short-term responsiveness, while DCEE quantifies long-term impacts, and together they can inform outcome selection, intervention design, and iterative refinement. Moreover, they can help assess whether proximal outcomes serve as meaningful surrogates for distal goals and guide how short-term and long-term objectives might be balanced in both offline and real-time adaptive settings. We provide a detailed discussion of these practical and scientific considerations in Supplementary Material G.
Future work could extend our approach in several directions. First, our method assumes no unmeasured confounding and a correctly specified propensity score, and both assumptions are satisfied in MRTs but not necessarily in observational studies; incorporating doubly robust estimation and methods to address unmeasured confounding would improve applicability to observational studies. Second, the current linear model for the DCEE could be extended to data-adaptive models, especially if the set of moderators $S_{t}$ is large. Third, our approach estimates an average causal effect across all participants; future research could adapt it to estimate person-specific heterogeneous effects. Fourth, linking proximal and distal outcomes through mediation analysis could help clarify mechanisms driving long-term behavior change. Fifth, extending the DCEE to allow general reference policies and multiple-step excursions would further broaden the applicability of the method.
Finally, future work could explore policy learning approaches to identify optimal treatment sequences that maximize long-term benefit. However, the numerous decision points pose challenges, including an exponentially growing policy space, possible positivity violations, and instability of inverse probability weights. Solutions may build upon ideas from reinforcement learning and dynamic treatment regime literature, such as making structural assumptions (e.g., Markov or restricted history dependence), searching within parameterized policy classes (e.g., using policy gradient methods), and compressing history into lower-dimensional representations using domain knowledge.
Acknowledgement
The author thanks the anonymous referees, the Associate Editor, and the Editor for their conscientious efforts and constructive comments which improved the quality of this paper.
@@ Acknowledgement
Supplementary Material A contains the extensions and details for the DCEE examples in Section 2.2. Supplementary Material B proves the identification result Theorem 2.1. Supplementary Material C proves the weighted average form of $\beta^{\star}$ given in Remark 4. Supplementary Material D derives the estimating function $\phi(\beta,\mu)$ in Eq. 10. Supplementary Material E proves the asymptotic normality Theorem 3.1. Supplementary Material F details the numerical computation of true parameter values in simulation studies. Supplementary Material G discusses practical considerations for using CEE and DCEE to inform intervention design.
Data Availability
The HeartSteps I MRT data that support the findings of this paper are publicly available at https://github.com/klasnja/HeartStepsV1. The code to reproduce the simulation and data analysis results can are available at https://github.com/tqian/paper_DCEE/tree/v1.0. The proposed estimator for distal causal excursion effect (DCEE) is implemented in the dcee() function in R package MRTAnalysis (Qian et al., 2025).
@@
We provide the detailed derivation for the four DCEE Examples in Section 2.2, where because Assumption 1 holds in all these examples, we can invoke Theorem 2.1 and in particular Eq. 5 to calculate $\tau(t)$. We also present more general versions of Example 3 and Example 4 to illustrate additional nuanced effects that are captured by the DCEE.